# Layer-wise Update Aggregation with Recycling for Communication-Efficient Federated Learning

### v1

@@ Appendix A Appendix
A.1 Theoretical Analysis
@@
Herein, we analyze the convergence properties of FedAvg as follows. First, the following Lemma is a slightly refined version of Lemma 3 in Reddi et al. [2021]. This Lemma is also used as Lemma 2 in Yang et al. [2021].
Lemma A.1.
@@
Then, by unrolling the recursion until $k-1$ goes to 00, we have
@@
$\sum_{t=0}^{T-1}\mathbb{E}\left[\left\|n_{t}\right\|^{2}\right]\leq 4T\tau^{2}\sigma_{L}^{2}+8T\tau^{2}\sigma_{G}^{2}+8\kappa\tau^{2}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\|\nabla F(\mathbf{x}_{t})\right\|^{2}\right]+\frac{8\tau L^{2}}{m}\sum_{t=0}^{T-1}\sum_{i=1}^{m}\sum_{j=0}^{\tau-1}\mathbb{E}\left[\left\|\mathbf{x}_{t,j}^{i}-\mathbf{x}_{t}\right\|^{2}\right],$ (12)
@@
where $\kappa$ is the ratio of the recycled update norm to the full update norm. Because all the gradients at the layers not recycled are zeroed out, the ratio $\kappa$ lies between 0 and 1; $10<\kappa<$ 1. ∎
Lemma A.3.
@@
where (21) holds if $\eta\leq\frac{1}{\mathcal{L}\tau}$.
Summing up (21) across $T$ communication rounds, we have
@@
Based on Lemma 3.1, the right-hand side can be re-written as follows.
@@
Now, we can derive the following Theorem based on Lemma A.1, A.3, and 3.1 as follows.
Theorem A.4.
Under assumption 1$\sim$3, if the learning rate $\eta\leq\frac{1-16\kappa}{6\sqrt{30}\mathcal{L}\tau}$ and $\kappa<\frac{1}{16}$, we have
@@
Based on Lemma A.1 and A.3, we have
@@
A.2 Experimental Settings in details
Implementation Details – All our experiments are conducted on a GPU cluster which contains 2 NVIDIA A6000 GPUs per machine. We use TensorFlow 2.15.0 for training and MPI for model aggregations. All individual experiments are performed at least three times, and the average accuracies are reported. The total number of clients is 128 and randomly chosen 32 clients participate in every communication round. We use mini-batch SGD with momentum (0.9) as the local optimizer. Table 5 shows the hyper-parameter settings for all our experiments, used for not only our method but also for other SOTA methods.
@@ A.2 Experimental Settings in details
Artificial Data Heterogeneity – For benchmark datasets that are not naturally non-IID, we generate an artificial data distributions using Dirichlet’s distributions. To evaluate the performance of our proposed method under realistic FL environments, the concentration coefficient $\alpha$ is configured as $0.1$ for CIFAR-10, CIFAR-100, and FEMNIST, and as $0.5$ for AG News. Note that these small concentration coefficient values represent the highly heterogeneous numbers of local samples across clients as well as imbalance of the samples across the labels.
Algorithm-Specific Hyperparameter Selection – Here, we summarize the hyper-parameter settings used to reproduce other SOTA methods, primarily following the configurations outlined in the original papers. We find algorithm-specific hyper-parameters using a grid search that achieve accuracy reasonably close to the baseline algorithm (FedAvg) while minimizing communication costs, and then measure the validation accuracy as shown in Section 4.1. Table 6 and Table 7 show the hyper-parameter settings for SOTA methods and experiments shown in Table 3 and 4, respectively. When running FedPara, both convolution layers and fully connected (FC) layers are re-paraeterized using their proposed method. All hyperparameters shown in Table 6, 7 are defined in the original papers.
@@
FedDropoutAvg $fdr$ (federated dropout rate) 0.5 0.4 0.75 0.5
@@
A.3 Extra Results
Sensitivity on $\delta$ – We performed a grid search for each dataset to find the best $\delta$ setting which yield the reasonable accuracy together with the maximum communication cost reduction. Table 8$\sim$11 show the four benchmarks’ accuracy and communication costs corresponding to various $\delta$ settings.
Learning Curves – Figure 5,6, and 7 show the learning curve comparisons for CIFAR-100, FEMNIST, and AG News benchmarks, respectively. To highlight the difference clearly, we chose only 3 representative methods and compare their curves to that of FedLUAR. It is clearly shown that FedLUAR achieves virtually the same accuracy as FedAvg while having a significantly reduced communication cost. These results well prove the efficacy of the proposed update recycling method.
@@ A.3 Extra Results
Table 8: The CIFAR-10 (ResNet20) classification performance with varying $\delta$ settings. The accuracy is almost not reduced until $\delta$ increases to $12$.
@@
Table 9: The CIFAR-100 (WideResNet28) classification performance with varying $\delta$ settings. The accuracy remains nearly the same until $\delta$ grows to $12$.
@@
Table 10: The FEMNIST (CNN) classification performance with varying $\delta$ settings. The communication cost is remarkably reduced while maintaining the accuracy until $\delta$ increase to $2$.
@@
Table 11: The AG news (DistillBERT) classification performance with varying $\delta$ settings. The communication cost is remarkably reduced while maintaining the accuracy.
@@
Figure 7: The learning curve comparisons for AG News (DistillBERT). The x-axis represents the communication cost relative to FedAvg. Likely to other behcnamrks, FedLUAR achieves the best accuracy faster than other SOTA methods.

### v2

@@ Appendix A Appendix
The appendix is structured as follows:
  • Section A.1 briefly announce broader impacts of this study.
  • Section A.2 provides problem definition, assumptions, and proofs of our theoretical analysis.
  • Section A.3 presents detailed experimental settings.
  • Section A.4 presents additional experimental results and analyses.
A.1 Broader Impacts
We do not anticipate any negative societal impact from our research. The proposed method accelerates neural network training in the context of Federated Learning. Faster training implies that target model accuracy can be achieved with fewer training iterations. As a result, it contributes to lower power consumption and a reduced carbon footprint.
A.2 Theoretical Analysis
@@
Herein, we analyze the convergence properties of FedAvg as follows. First, the following Lemma is a slightly refined version of Lemma 3 in [30]. This Lemma is also used as Lemma 2 in [38].
Lemma 3.
@@
Then, by unrolling the recursion until k1k-1 goes to 0, we have
@@
$\sum_{t=0}^{T-1}\mathbb{E}\left[\left\|n_{t}\right\|^{2}\right]$ $\leq 4T\tau^{2}\sigma_{L}^{2}+8T\tau^{2}\sigma_{G}^{2}+8\kappa\tau^{2}\sum_{t=0}^{T-1}\mathbb{E}\left[\left\|\nabla F(\mathbf{x}_{t})\right\|^{2}\right]$
$\qquad\qquad+\frac{8\tau L^{2}}{m}\sum_{t=0}^{T-1}\sum_{i=1}^{m}\sum_{j=0}^{\tau-1}\mathbb{E}\left[\left\|\mathbf{x}_{t,j}^{i}-\mathbf{x}_{t}\right\|^{2}\right],$ (12)
@@
where $\kappa$ is the ratio of the recycled update norm to the full update norm. Because all the gradients at the layers not recycled are zeroed out, the ratio $\kappa$ lies between 0 and 1; $0<\kappa<1$. ∎
Lemma 5.
@@
where (21) holds if $\eta\leq\frac{1}{\mathcal{L}\tau}$. Summing up (21) across $T$ communication rounds, we have
@@
Based on Lemma 1, the right-hand side can be re-written as follows.
@@
Now, we can derive the following Theorem based on Lemma 3, 5, and 1 as follows.
Theorem 6.
[…]
@@
Based on Lemma 3 and 5, we have
@@
A.3 Experimental Settings in details
Implementation Details – All our experiments are conducted on a GPU cluster that contains 2 NVIDIA A6000 GPUs per machine. We use TensorFlow 2.15.0 for training and MPI for model aggregations. All individual experiments are performed at least three times, and the average accuracies are reported. The total number of clients is 128 and randomly chosen 32 clients participate in every communication round. We use mini-batch SGD with momentum (0.9) as the local optimizer. Table 6 shows the hyper-parameter settings for all our experiments, used not only for our method but also for other SOTA methods.
@@ A.3 Experimental Settings in details
Artificial Data Heterogeneity – For benchmark datasets that are not naturally non-IID, we generate artificial data distributions using Dirichlet’s distributions. To evaluate the performance of our proposed method under realistic FL environments, the concentration coefficient $\alpha$ is configured as $0.1$ for CIFAR-10, CIFAR-100, and FEMNIST, and as $0.5$ for AG News. Note that these small concentration coefficient values represent highly heterogeneous distributions of local samples across clients as well as imbalance in the number of samples across labels.
Algorithm-Specific Hyperparameter Selection – Here, we summarize the hyper-parameter settings used to reproduce other SOTA methods, primarily following the configurations outlined in the original papers. We find algorithm-specific hyper-parameters using a grid search that achieve accuracy reasonably close to the baseline algorithm (FedAvg) while minimizing communication costs, and then measure the validation accuracy as shown in Section 4.1. Table 7 and Table 8 show the hyper-parameter settings for SOTA methods and experiments shown in Table 2 and 3, respectively. When running FedPara, both convolution layers and fully connected (FC) layers are re-parameterized using their proposed method. All hyperparameters shown in Table 7 and 8 are defined in the original papers.
@@
FedDropoutAvg fdrfdr (federated dropout rate) 0.5 0.4 0.75 0.5
FedBAT $\rho,\phi$ (coefficient, warm-up ratio) 6, 0.5 6, 0.5 6, 0.5 6, 0.5
@@
A.4 Extra Results
Sensitivity on $\delta$ – We performed a grid search for each dataset to find the best $\delta$ setting which yields reasonable accuracy together with the maximum communication cost reduction. Table 912 show the four benchmarks’ accuracy and communication costs corresponding to various $\delta$ settings.
How much could be recycled safely? – We also investigate the impact of $\delta$ on model accuracy and communication costs. We divide the number of model aggregations at each layer by the total number of communication rounds to calculate the layer-wise communication cost. Then, we sum up the calculated layer-wise costs to get the total communication cost. Intuitively, the larger the $\delta$, the lower the communication cost. However, the model accuracy is expected to drop as more layers have their updates recycled.
Table 9 and 10 show CIFAR-10 and CIFAR-100 experimental results for various $\delta$ settings. One key observation is that the accuracy is almost not degraded when LUAR is applied with $\delta\leq 12$ for both datasets. This means that many network layers have quite stable gradient dynamics, and thus their updates can be safely recycled. In addition, the accuracy is hardly reduced until the communication cost is reduced by almost $50\%$. This is a significant benefit especially in FL environments where the network bandwidth is extremely limited.
Results under varying degrees of data heterogeneity – The degree of non-IIDness strongly affects the training efficiency of Federated Learning methods. We conducted additional experiments to demonstrate that FedLUAR is robust to various degrees of non-IIDness. Table 13 and Table 14 show CIFAR-10 and AG News experimental results for various Dirchlet concentration factor $\alpha$ settings. In both benchmarks, FedLUAR achieves comparable accuracy to FedAvg regardless of $\alpha$, while considerably reducing the communication cost. Therefore, we conclude that FedLUAR is robust to the degree of non-IIDness.
Performance under different numbers of active clients – We conducted additional ablation study using different numbers of active clients. Table 15 and Table 16 show that, regardless of the total number of clients, FedLUAR achieves accuracy comparable to FedAvg while significantly reducing the communication cost. This ablation study demonstrates the superior scalability of FedLUAR.
Learning Curves – Figure 5 and 6 show the learning curve comparisons for CIFAR-100 and FEMNIST benchmarks, respectively. To highlight the difference clearly, we chose only 3 representative methods and compare their curves to those of FedLUAR. It is clearly shown that FedLUAR achieves virtually the same accuracy as FedAvg while having a significantly reduced communication cost. These results well prove the efficacy of the proposed update recycling method.
@@ A.4 Extra Results
Table 9: The CIFAR-10 (ResNet20) classification performance with varying $\delta$ settings.
@@
Table 10: The CIFAR-100 (WideResNet28) classification performance with varying $\delta$ settings.
@@
Table 11: The FEMNIST (CNN) classification performance with varying $\delta$ settings.
@@
Table 12: The AG news (DistillBERT) classification performance with varying $\delta$ settings.
Method $\alpha=0.1$ $\alpha=0.5$ $\alpha=1.0$
Comm Acc Acc Acc
FedAvg $1.00$ $61.27\%$ $76.54\%$ $79.73\%$
FedLUAR $0.47$ $60.15\%$ $76.04\%$ $79.50\%$
Table 13: CIFAR-10 with various Dirchlet concentration factor $\alpha$ settings. The number of recycled layers, $\delta=10$ out of 20 layers in ResNet20.
Method $\alpha=0.1$ $\alpha=0.5$ $\alpha=1.0$
Comm Acc Acc Acc
FedAvg $1.00$ $81.53\%$ $82.66\%$ $83.22\%$
FedLUAR $0.17$ $81.88\%$ $82.80\%$ $82.75\%$
Table 14: AG News with various Dirchlet concentration factor $\alpha$ settings. The number of recycled layers, $\delta=30$ out of 40 layers in DistilBERT.
Method 64 (0.5) 128 (0.25) 256 (0.125)
Comm Acc Acc Acc
FedAvg $1.00$ $54.24\%$ $61.27\%$ $57.81\%$
FedLUAR $0.48$ $53.34\%$ $60.15\%$ $57.81\%$
Table 15: The $\delta=10$ out of 20 layers in ResNet20 for FedLUAR. 64, 128, and 256 indicate the total number of clients, and the numbers in parentheses (0.5, 0.25, 0.125) represent the client activation ratio on CIFAR-10.
Method 64 (0.5) 128 (0.25) 256 (0.125)
Comm Acc Acc Acc
FedAvg $1.00$ $66.31\%$ $71.01\%$ $75.97\%$
FedLUAR $0.14$ $68.15\%$ $73.17\%$ $76.72\%$
Table 16: The $\delta=2$ out of 4 layers in CNN for FedLUAR. 64, 128, and 256 indicate the total number of clients, and the numbers in parentheses (0.5, 0.25, 0.125) represent the client activation ratio on FEMNIST.